OpenStudy (anonymous):

49x(square)-7/3x+1/36=0 what can be x? using a(square)-2ab+b(square)

OpenStudy (anonymous):

\[49x^2-\frac{7}{3}x+\frac{1}{36}\] is a perfect square. it is \[(7x-\frac{1}{6})^2\]